OpenStudy (andriani):

P(k) 1^3 + 2^3 + 3^3 + ..... + k^3 = (1+2+3+...+k)^2 P(k+1) 1^3 + 2^3 + 3^3 + .... + k^3 +(k+1)^3 = (1 + 2 + 3 + .... + k(k+1))^2 Please help me to prove that it's true

jhonyy9 (jhonyy9):

what method of proof you wan using for this ?

OpenStudy (andriani):

Math induction. It's on the second step

Nnesha (nnesha):

i believe you already did the first step where we assume the statement is true for n=1 \[ \rm \color{Red}{1^3 + 2^3 + 3^3 + ..... + k^3} = \color{blue}{(1+2+3+...+k)^2} \] \[\large\rm \color{Red}{1^3 + 2^3 + 3^3 + .... + k^3 }+(k+1)^3 = (1 + 2 + 3 + .... + k(k+1))^2 \] can we replace red part with blue one right ?? \[\large\rm \color{blue}{(1+2+3+...+k)^2}+(k+1)^3 = (1 + 2 + 3 + .... + k(k+1))^2 \]

Nnesha (nnesha):

wait can i see the original question ?? :P

OpenStudy (andriani):

1^3 + 2^3 + 3^3 + ... + n^3 = (1 + 2 + 3 + ... + n) ^2 Thank u soo muchhh